# Moving toward best practice when using propensity score weighting in survey observational studies

### v1

@@ 3.1 Augmented weighting estimating equations for survey observational data
In the survey setting, the augmented weighting estimator $\widehat{\tau}_{h}^{\text{AW}}$ remains consistent for $\tau_{h}$ provided that the population-level propensity score model is correctly specified, irrespective of the accuracy of the outcome models. When the population-level propensity score model $e_{\text{sp}}(\mathbf{X})$ is misspecified but the outcome models are accurate, this estimator consistently estimates $\widetilde{\tau}_{h}$, which only equals to $\tau_{h}$ when the tilting function $h(\mathbf{X})$ does not depend on $e_{\text{sp}}(\mathbf{X})$.
In the following, we introduce three variants of the augmented weighting estimators that are applicable to the survey observational data, all connected to and motivated from Equation (11).
@@ 3.2 One-step moment estimator
The first augmented estimator is the moment estimator (MOM). Under the balancing weights framework, the moment estimator can be adapted for any specified target population defined by covariate density $g(\mathbf{X})$ or tilting function $h(\mathbf{X})$.[27] We now extend this traditional augmented estimator to the survey setting by augmenting the propensity score weighting estimator in Equation (3) with the outcome regression model. We denote the regression model as the conditional expectation of the outcome given covariates over treatment and control group in the observed sample as $m_{1}^{\text{MOM}}(\mathbf{X};\mathbf{\alpha}_{1})$ and $m_{0}^{\text{MOM}}(\mathbf{X};\mathbf{\alpha}_{0})$. Here, $\mathbf{\alpha}_{1}$ and $\mathbf{\alpha}_{0}$ are vectors of parameters estimated from the observed data in each arm, representing the relationship between covariates $\mathbf{X}$ and the expected outcome in the treatment and control groups, respectively. By plugging in $m_{1}^{\text{MOM}}(\mathbf{X};\widehat{\mathbf{\alpha}}_{1})$ and $m_{0}^{\text{MOM}}(\mathbf{X};\widehat{\mathbf{\alpha}}_{0})$ into $\widehat{\tau}_{h}^{\text{AW}}$, the moment estimator can be subsequently expressed as,
$\widehat{\tau}_{h}^{\text{MOM}}$ $=\frac{\sum_{i=1}^{n}h(\mathbf{X}_{i})/P(S_{i}=1\mid\mathbf{X}_{i})\left[m_{1}^{\text{MOM}}(\mathbf{X}_{i};\widehat{\mathbf{\alpha}}_{1})-m_{0}^{\text{MOM}}(\mathbf{X}_{i};\widehat{\mathbf{\alpha}}_{0})\right]}{\sum_{i=1}^{n}h(\mathbf{X}_{i})/P(S_{i}=1\mid\mathbf{X}_{i})}$ (13)
@@
$\quad+\frac{\sum_{i=1}^{n}\omega_{1,p}(\mathbf{X}_{i})Z_{i}\left[Y_{i}-m_{1}^{\text{MOM}}(\mathbf{X}_{i};\widehat{\mathbf{\alpha}}_{1})\right]}{\sum_{i=1}^{n}\omega_{1,p}(\mathbf{X}_{i})Z_{i}}-\frac{\sum_{i=1}^{n}\omega_{0,p}(\mathbf{X}_{i})(1-Z_{i})\left[Y_{i}-m_{0}^{\text{MOM}}(\mathbf{X}_{i};\widehat{\mathbf{\alpha}}_{0})\right]}{\sum_{i=1}^{n}\omega_{0,p}(\mathbf{X}_{i})(1-Z_{i})}$
In this moment estimator, survey weights must be incorporated in three different places. First, survey weights are used to estimate the population-level propensity score, $\widehat{e}_{\text{sp}}(\mathbf{X}_{i};\widehat{\boldsymbol{\beta}})$, via weighted logistics regression. Second, the estimated $\widehat{e}_{\text{sp}}(\mathbf{X}_{i};\widehat{\boldsymbol{\beta}})$ is utilized to construct the tilting function $h(\mathbf{X}_{i})=h(\widehat{e}_{\text{sp}}(\mathbf{X}_{i};\widehat{\boldsymbol{\beta}}))$, which defines the target population estimand, as illustrated in Table 1. Third, survey weights are further incorporated into the construction of the balancing weight pair, $\omega_{1,p}(\mathbf{X}_{i})=h(\widehat{e}_{\text{sp}}(\mathbf{X}_{i};\widehat{\boldsymbol{\beta}}))/(p_{1}(\mathbf{X}_{i})\widehat{e}_{\text{sp}}(\mathbf{X}_{i};\widehat{\boldsymbol{\beta}}))$ and $\omega_{0,p}(\mathbf{X}_{i})=h(\widehat{e}_{\text{sp}}(\mathbf{X}_{i};\widehat{\boldsymbol{\beta}})/(p_{0}(\mathbf{X}_{i})(1-\widehat{e}_{\text{sp}}(\mathbf{X}_{i};\widehat{\boldsymbol{\beta}})))$, effectively reweighting the sample to represent the target population through the tilting function $h(\widehat{e}_{\text{sp}}(\mathbf{X}_{i};\widehat{\boldsymbol{\beta}})$ and additional multiplication with the sampling probabilities $p_{z}(\mathbf{X}_{i})$. Here, the conditional expectations, $m_{1}^{\text{MOM}}(\mathbf{X}_{i};\mathbf{\alpha}_{1})$ and $m_{0}^{\text{MOM}}(\mathbf{X}_{i};\mathbf{\alpha}_{0})$, are regression models applied within the sample to recover the missing potential outcomes. Detailed proof for the consistency of $\widehat{\tau}_{h}^{\text{MOM}}$ is in Web Appendix A.5. Equation 13 can be decomposed into two components: $\widehat{\tau}_{h}^{\text{MOM}}=\widehat{\mu}_{1}^{\text{MOM}}-\widehat{\mu}_{0}^{\text{MOM}}$, where
$\widehat{\mu}_{1}^{\text{MOM}}$ $=\frac{\sum_{i=1}^{n}h(\mathbf{X}_{i})/P(S_{i}=1\mid\mathbf{X}_{i})m_{1}^{\text{MOM}}(\mathbf{X}_{i};\widehat{\mathbf{\alpha}}_{1})}{\sum_{i=1}^{n}h(\mathbf{X}_{i})/P(S_{i}=1\mid\mathbf{X}_{i})}+\frac{\sum_{i=1}^{n}\omega_{1,p}(\mathbf{X}_{i})Z_{i}\left[Y_{i}-m_{1}^{\text{MOM}}(\mathbf{X}_{i};\widehat{\mathbf{\alpha}}_{1})\right]}{\sum_{i=1}^{n}\omega_{1,p}(\mathbf{X}_{i})Z_{i}},$ (14)
and $\widehat{\mu}_{0}^{\text{MOM}}$ is the counterpart for the control group. Each of $\widehat{\mu}_{1}^{\text{MOM}}$ and $\widehat{\mu}_{0}^{\text{MOM}}$ is characterized by a regression model enhanced by weighted residuals for each group. Regarding the augmentation properties of $\widehat{\tau}_{h}^{\text{MOM}}$, consider the expression for $\widehat{\mu}_{1}^{\text{MOM}}$, which has the following asymptotic approximation,
@@
The second term in this equation vanishes when either the postulated regression model $m_{1}^{\text{MOM}}(\mathbf{X};\mathbf{\alpha}_{1})$ or the survey-weighted propensity score model $e_{\text{sp}}(\mathbf{X};\boldsymbol{\beta})$ is correctly specified. When the tilting function $h(\mathbf{X})$ is specified as a constant proportional function, such as $h(\mathbf{X})\propto 1$, it corresponds to inverse probability weighting and defines the entire population as the target population. The estimand in this case is the population average treatment effect (PATE). Under this configuration, this moment estimator $\widehat{\tau}_{h}^{\text{MOM}}$ exhibits double robustness, meaning that consistency is guaranteed if either the outcome model $m_{1}^{\text{MOM}}(\mathbf{X};\mathbf{\alpha}_{1})$ or the propensity score model $e_{\text{sp}}(\mathbf{X};\boldsymbol{\beta})$ is correctly specified. Similarly, when $h(\mathbf{X})$ corresponds to trimming functions, it depends directly on $\mathbf{X}$ and does not rely on the correct specification of $e_{\text{sp}}(\mathbf{X};\boldsymbol{\beta})$. In such cases, $\widehat{\tau}_{h}^{\text{MOM}}$ retains double robustness.
In more general cases, when the tilting function $h(\mathbf{X})$ explicitly depends on $e_{\text{sp}}(\mathbf{X})$, as illustrated in Table 1, the estimator $\widehat{\mu}_{1}^{\text{MOM}}$ exhibits only single robustness. In this context, consistency is ensured only if the propensity score model $e_{\text{sp}}(\mathbf{X};\boldsymbol{\beta})$ is correctly specified, which allows for unbiased estimation of the first term in Equation 15. However, as demonstrated by Mao et al.,[27] integrating an outcome model into the weighting estimator can improve efficiency compared to weighting alone. Even in cases where $e_{\text{sp}}(\mathbf{X};\boldsymbol{\beta})$ is misspecified, leveraging $m_{1}^{\text{MOM}}(\mathbf{X};\mathbf{\alpha}_{1})$ can help reduce residual variance, thereby enhancing the estimator’s precision.
@@ 3.3 Clever covariate regression estimator
An alternative strategy for constructing an augmented estimator involves incorporating functions of estimated propensity scores into the outcome regression model as an additional covariate. Scharfstein et al.[29] introduced a specific version of this approach by adding the inverse propensity score into outcome models. Later, Bang et al.[25] demonstrated that this estimator is essentially a solution to the augmented weighting estimating equation, making the inclusion of a single propensity-score-based covariate sufficient to achieve double robustness for augmented IPW estimators. This method is often referred to as the “clever covariate” estimator (CVR) because the additional covariate “cleverly” adjusts for potential misspecification of the outcome model by introducing a coarse but sufficient representation of all covariates, thereby ensuring robustness. Moreover, as noted by Robins et al., [26] when the CVR is structured with an identity link, it aligns with the targeted maximum likelihood estimation (TMLE) for marginal means over the target population, using, in our case, population-level survey-weighted balancing weights as covariates to re-target and correct the original naive outcome model to better estimate potential outcomes in the target population.
Under the context of survey observational data, by solving the augmented weighting estimating equation3.1 and replacing the conventional outcome models $m_{1}(\mathbf{X}_{i})$ and $m_{0}(\mathbf{X}_{i})$ with generalized linear regression models $m_{1}^{\text{CVR}}\left(\mathbf{X};\widehat{\boldsymbol{\alpha}}_{1}^{\text{CVR}}\right)=g^{-1}\{m_{1}(\mathbf{X};\widehat{\alpha}_{1})+\widehat{\phi}_{1}\omega_{1,p}(\mathbf{X})\}$ and $m_{0}^{\text{CVR}}\left(\mathbf{X};\widehat{\boldsymbol{\alpha}}_{0}^{\text{CVR}}\right)=g^{-1}\{m_{0}(\mathbf{X};\widehat{\alpha}_{0})+\widehat{\phi}_{0}\omega_{0,p}(\mathbf{X})\}$, where $g$ is the canonical link function, we can construct the clever covariate estimator adapted to survey observational data as follows:
$\quad\widehat{\tau}_{h}^{\text{CVR}}$
$=\widehat{\tau}_{1}^{\text{CVR}}-\widehat{\tau}_{0}^{\text{CVR}}$ (16)
$=\frac{\sum_{i=1}^{n}\displaystyle h(\mathbf{X}_{i})/p(S_{i}=1\mid\mathbf{X}_{i})m_{1}^{\text{CVR}}\left(\mathbf{X}_{i};\widehat{\boldsymbol{\alpha}}_{1}^{\text{CVR}}\right)}{\sum_{i=1}^{n}h(\mathbf{X}_{i})/p(S_{i}=1\mid\mathbf{X}_{i})}+\frac{\sum_{i=1}^{n}\omega_{1,p}(\mathbf{X}_{i},\widehat{\boldsymbol{\beta}})Z_{i}\left(Y_{i}-m_{1}^{\text{CVR}}\left(\mathbf{X}_{i};\widehat{\boldsymbol{\alpha}}_{1}^{\text{CVR}}\right)\right)}{\sum_{i=1}^{n}\omega_{1,p}\left(\mathbf{X}_{i};\widehat{\boldsymbol{\beta}}\right)Z_{i}}$
$-\frac{\sum_{i=1}^{n}h(\mathbf{X}_{i})/p(S_{i}=1\mid\mathbf{X}_{i})m_{0}^{\text{CVR}}\left(\mathbf{X}_{i};\widehat{\boldsymbol{\alpha}}_{0}^{\text{CVR}}\right)}{\sum_{i=1}^{n}h(\mathbf{X}_{i})/p(S_{i}=1\mid\mathbf{X}_{i})}-\frac{\sum_{i=1}^{n}\omega_{0,p}(\mathbf{X}_{i},\widehat{\boldsymbol{\beta}})(1-Z_{i})\left(Y_{i}-m_{0}^{\text{CVR}}\left(\mathbf{X}_{i};\widehat{\boldsymbol{\alpha}}_{0}^{\text{CVR}}\right)\right)}{\sum_{i=1}^{n}\omega_{0,p}\left(\mathbf{X}_{i};\widehat{\boldsymbol{\beta}}\right)(1-Z_{i})}$
@@
$\widehat{\tau}_{h}^{\text{CVR}}=\frac{\sum_{i=1}^{n}\displaystyle h(\mathbf{X}_{i})/p(S_{i}=1\mid\mathbf{X}_{i})\left(m_{1}^{\text{CVR}}\left(\mathbf{X}_{i};\widehat{\boldsymbol{\alpha}}_{1}^{\text{CVR}}\right)-m_{0}^{\text{CVR}}\left(\mathbf{X}_{i};\widehat{\boldsymbol{\alpha}}_{0}^{\text{CVR}}\right)\right)}{\sum_{i=1}^{n}h(\mathbf{X}_{i})/p(S_{i}=1\mid\mathbf{X}_{i})}$ (17)
Here, by substituting the expression of balancing weights pair $(\omega_{1,p}(\mathbf{X}_{i}),\omega_{0,p}(\mathbf{X}_{i}))$ as the clever covariates for each arm, we have
$m_{1}^{\text{CVR}}\left(\mathbf{X}_{i};\widehat{\boldsymbol{\alpha}}_{1}^{\text{CVR}}\right)=g^{-1}\left\{m_{1}(\mathbf{X}_{i};\widehat{\alpha}_{1})+\widehat{\phi}_{1}h(\mathbf{X}_{i})\,/\,\big{(}P(S_{i}=1\mid Z_{i}=1,\mathbf{X}_{i})e_{\text{sp}}(\mathbf{X}_{i};\widehat{\boldsymbol{\beta}})\big{)}\right\}$
and $m_{0}^{\text{CVR}}\left(\mathbf{X}_{i};\widehat{\boldsymbol{\alpha}}_{0}^{\text{CVR}}\right)=g^{-1}\left\{m_{0}(\mathbf{X}_{i};\widehat{\alpha}_{0})+\widehat{\phi}_{0}h(\mathbf{X}_{i})\,/\,\big{(}P(S_{i}=1\mid Z_{i}=0,\mathbf{X}_{i})(1-e_{\text{sp}}(\mathbf{X}_{i};\widehat{\boldsymbol{\beta}}))\big{)}\right\}.$ In this estimator, survey weights need to be incorporated three times in the clever covariate method: during survey-weighted propensity score regression for $e_{\text{sp}}(\mathbf{X}_{i};\widehat{\boldsymbol{\beta}})$ , when formulating balancing weights as clever covariates, and in standardizing the outcome predictions toward the target estimands.
The clever covariate estimator is asymptotically equivalent to the moment estimator in that both are solutions to the augmented weighting estimating equation and share the same influence function. When the tilting function $h(\mathbf{X})\propto 1$, the clever covariate method becomes equivalent to the augmented IPW estimator, which exhibits double robustness, ensuring consistency if either the outcome model (without clever covariates) $m_{z}\left(\mathbf{X};\widehat{\boldsymbol{\alpha}}_{z}\right)$ or the propensity score model $e_{\text{sp}}(\mathbf{X};\widehat{\boldsymbol{\beta}})$ is correctly specified. In more general cases, when $h(\mathbf{X})$ explicitly depends on $e_{\text{sp}}(\mathbf{X})$, it may improve efficiency over simple weighting estimators discussed in Section 2. However, this requires the correct specification of $e_{\text{sp}}(\mathbf{X};\widehat{\boldsymbol{\beta}})$; otherwise, the estimator converges to a biased weighted average treatment effect estimand $\widetilde{\tau}_{h}$, where the weights depend on the misspecified tilting function $h(e_{\text{sp}}(\mathbf{X};\widetilde{\boldsymbol{\beta}}))$.
@@
In the absence of survey sampling, Kang and Schafer [30] pointed out that, when the outcome models $m_{1}\left(\mathbf{X};\widehat{\boldsymbol{\alpha}}_{1}\right)$ and $m_{0}\left(\mathbf{X};\widehat{\boldsymbol{\alpha}}_{0}\right)$ are already correctly specified, the clever covariates as coarse summaries of the covariates may lead to slight overfitting by adding mean-zero noise through weighted residuals. When the survey-weighted propensity score model is correctly specified, the clever covariate method ensures consistent estimation of target parameters, provided that the conditional mean of the outcome varies smoothly with both the survey-weighted propensity score and the sampling probabilities. In theory, this smooth variation remains essential for maintaining the consistency and robustness of the clever covariate method, even in complex survey designs. However, incorporating survey weights twice in constructing the clever covariates may potentially exacerbate the extrapolation problem noted by Robins et al.[26] because the inflation of extreme value can lead to finite-sample bias, making the clever covariate estimator less stable compared to the weighted regression methods introduced in Section 3.4.

### v2

@@ 3.1 Augmented weighting estimating equations for survey observational data
The augmented weighting estimator $\widehat{\tau}_{h}^{\text{AW}}$ remains consistent for $\tau_{h}$ provided that the population-level propensity score model is correctly specified, irrespective of the accuracy of the outcome models. When the population-level propensity score model $e_{\text{sp}}(\mathbf{X})$ is misspecified but the outcome models are correct, this estimator consistently estimates $\widetilde{\tau}_{h}$.
[…]
In the following, we introduce three variants of the augmented weighting estimators that are applicable to survey observational data, all connected to Equation (13).
@@ 3.2 One-step moment estimator
The first augmented estimator is the one-step moment estimator (MOM). Under the balancing weights framework, the moment estimator can be adapted for any specified target population defined by $h(\mathbf{X})$.[21] We denote the regression model as the conditional expectation of the outcome given covariates over treatment and control group in the observed sample as $m_{1}^{\text{MOM}}(\mathbf{X};\mathbf{\alpha}_{1})$ and $m_{0}^{\text{MOM}}(\mathbf{X};\mathbf{\alpha}_{0})$, with $\mathbf{\alpha}_{1}$ and $\mathbf{\alpha}_{0}$ as vectors of regression parameters. By plugging in $m_{1}^{\text{MOM}}(\mathbf{X};\widehat{\mathbf{\alpha}}_{1})$ and $m_{0}^{\text{MOM}}(\mathbf{X};\widehat{\mathbf{\alpha}}_{0})$ into $\widehat{\tau}_{h}^{\text{AW}}$, the moment estimator can be expressed as,
$\widehat{\tau}_{h}^{\text{MOM}}$ $=\frac{\sum_{i=1}^{n}h(\mathbf{X}_{i})/p(\mathbf{X}_{i})\left[m_{1}^{\text{MOM}}(\mathbf{X}_{i};\widehat{\mathbf{\alpha}}_{1})-m_{0}^{\text{MOM}}(\mathbf{X}_{i};\widehat{\mathbf{\alpha}}_{0})\right]}{\sum_{i=1}^{n}h(\mathbf{X}_{i})/p(\mathbf{X}_{i})}$ (14)
@@
$\quad+\frac{\sum_{i=1}^{n}\omega_{1,p}(\mathbf{X}_{i})Z_{i}\left[Y_{i}-m_{1}^{\text{MOM}}(\mathbf{X}_{i};\widehat{\mathbf{\alpha}}_{1})\right]}{\sum_{i=1}^{n}\omega_{1,p}(\mathbf{X}_{i})Z_{i}}-\frac{\sum_{i=1}^{n}\omega_{0,p}(\mathbf{X}_{i})(1-Z_{i})\left[Y_{i}-m_{0}^{\text{MOM}}(\mathbf{X}_{i};\widehat{\mathbf{\alpha}}_{0})\right]}{\sum_{i=1}^{n}\omega_{0,p}(\mathbf{X}_{i})(1-Z_{i})}.$
@@
Upon closer inspection, Equation (14) can be decomposed into two components: $\widehat{\tau}_{h}^{\text{MOM}}=\widehat{\mu}_{1}^{\text{MOM}}-\widehat{\mu}_{0}^{\text{MOM}}$, where
$\widehat{\mu}_{1}^{\text{MOM}}$ $=\frac{\sum_{i=1}^{n}h(\mathbf{X}_{i})/p(\mathbf{X}_{i})m_{1}^{\text{MOM}}(\mathbf{X}_{i};\widehat{\mathbf{\alpha}}_{1})}{\sum_{i=1}^{n}h(\mathbf{X}_{i})/p(\mathbf{X}_{i})}+\frac{\sum_{i=1}^{n}\omega_{1,p}(\mathbf{X}_{i})Z_{i}\left[Y_{i}-m_{1}^{\text{MOM}}(\mathbf{X}_{i};\widehat{\mathbf{\alpha}}_{1})\right]}{\sum_{i=1}^{n}\omega_{1,p}(\mathbf{X}_{i})Z_{i}},$ (15)
and $\widehat{\mu}_{0}^{\text{MOM}}$ is the counterpart for the control group. Each of $\widehat{\mu}_{1}^{\text{MOM}}$ and $\widehat{\mu}_{0}^{\text{MOM}}$ is characterized by a regression model enhanced by weighted residuals for each group. Regarding the augmentation properties of $\widehat{\tau}_{h}^{\text{MOM}}$, consider the expression for $\widehat{\mu}_{1}^{\text{MOM}}$, which converges in probability to
$\widehat{\mu}_{1}^{\text{MOM}}\xrightarrow{p}\frac{E\left[h(\mathbf{X})Y(1)\right]}{E\left[h(\mathbf{X})\right]}+\frac{1}{E\left[h(\mathbf{X})\right]}E\left[h(\mathbf{X})\left(\frac{Z}{e_{\text{sp}}(\mathbf{X})}-1\right)\left(Y(1)-m_{1}^{\text{MOM}}(\mathbf{X};\mathbf{\alpha}_{1})\right)\right].$ (16)
The second term in this equation vanishes when either the postulated regression model $m_{1}^{\text{MOM}}(\mathbf{X};\mathbf{\alpha}_{1})$ or the survey-weighted propensity score model $e_{\text{sp}}(\mathbf{X};\boldsymbol{\beta})$ is correctly specified. When the tilting function $h(\mathbf{X})$ is specified as a constant proportional function, such as $h(\mathbf{X})\propto 1$, it corresponds to IPW and defines the entire population as the target population (i.e., PATE is the estimand). Under this configuration, $\widehat{\tau}_{h}^{\text{MOM}}$ exhibits double robustness, meaning that consistency is guaranteed if either the outcome model $m_{1}^{\text{MOM}}(\mathbf{X};\mathbf{\alpha}_{1})$ or the propensity score model $e_{\text{sp}}(\mathbf{X};\boldsymbol{\beta})$ is correctly specified.
In more general cases, when the tilting function $h(\mathbf{X})$ explicitly depends on $e_{\text{sp}}(\mathbf{X})$, as illustrated in Table 1, the estimator $\widehat{\mu}_{1}^{\text{MOM}}$ is only singly robust. In this context, consistency is ensured only if the propensity score model $e_{\text{sp}}(\mathbf{X};\boldsymbol{\beta})$ is correctly specified, which allows for unbiased estimation of the first term in Equation (16). However, as demonstrated by Mao et al.,[21] integrating an outcome model into the weighting estimator can improve efficiency compared to weighting alone. Through simulation studies, we demonstrate in Section 4 that this observation continues to hold in survey observational studies.
@@ 3.3 Clever covariate regression estimator
An alternative strategy for constructing an augmented estimator involves incorporating functions of estimated propensity scores into the outcome regression model as an additional covariate. Scharfstein et al.[24] introduced a specific version of this approach by adding the inverse propensity score into outcome models. Later, Bang et al.[25] demonstrated that this estimator is essentially a solution to the augmented weighting estimating equation; hence in the context of IPW, the resulting estimator is also doubly robust. This method is also referred to as the “clever covariate” estimator (CVR), and we present a version of this estimator for survey observational studies. Specifically, we first fit the following modified outcome models $m_{1}(\mathbf{X})$ and $m_{0}(\mathbf{X})$ with the additional clever covariate:
[…]
where $g$ is the canonical link function. The clever covariate estimator is obtained by using the modified regression functions
$\widehat{\tau}_{h}^{\text{CVR}}$ $=\widehat{\tau}_{1}^{\text{CVR}}-\widehat{\tau}_{0}^{\text{CVR}}$ (17)
$=\frac{\sum_{i=1}^{n}\displaystyle h(\mathbf{X}_{i})/p(\mathbf{X}_{i})m_{1}^{\text{CVR}}\left(\mathbf{X}_{i};\widehat{\boldsymbol{\alpha}}_{1}^{\text{CVR}}\right)}{\sum_{i=1}^{n}h(\mathbf{X}_{i})/p(\mathbf{X}_{i})}+\frac{\sum_{i=1}^{n}\omega_{1,p}(\mathbf{X}_{i},\widehat{\boldsymbol{\beta}})Z_{i}\left(Y_{i}-m_{1}^{\text{CVR}}\left(\mathbf{X}_{i};\widehat{\boldsymbol{\alpha}}_{1}^{\text{CVR}}\right)\right)}{\sum_{i=1}^{n}\omega_{1,p}\left(\mathbf{X}_{i};\widehat{\boldsymbol{\beta}}\right)Z_{i}}$
$\quad-\frac{\sum_{i=1}^{n}h(\mathbf{X}_{i})/p(\mathbf{X}_{i})m_{0}^{\text{CVR}}\left(\mathbf{X}_{i};\widehat{\boldsymbol{\alpha}}_{0}^{\text{CVR}}\right)}{\sum_{i=1}^{n}h(\mathbf{X}_{i})/p(\mathbf{X}_{i})}-\frac{\sum_{i=1}^{n}\omega_{0,p}(\mathbf{X}_{i},\widehat{\boldsymbol{\beta}})(1-Z_{i})\left(Y_{i}-m_{0}^{\text{CVR}}\left(\mathbf{X}_{i};\widehat{\boldsymbol{\alpha}}_{0}^{\text{CVR}}\right)\right)}{\sum_{i=1}^{n}\omega_{0,p}\left(\mathbf{X}_{i};\widehat{\boldsymbol{\beta}}\right)(1-Z_{i})}.$
@@
$\widehat{\tau}_{h}^{\text{CVR}}=\frac{\sum_{i=1}^{n}\displaystyle h(\mathbf{X}_{i})/p(\mathbf{X}_{i})\left(m_{1}^{\text{CVR}}\left(\mathbf{X}_{i};\widehat{\boldsymbol{\alpha}}_{1}^{\text{CVR}}\right)-m_{0}^{\text{CVR}}\left(\mathbf{X}_{i};\widehat{\boldsymbol{\alpha}}_{0}^{\text{CVR}}\right)\right)}{\sum_{i=1}^{n}h(\mathbf{X}_{i})/p(\mathbf{X}_{i})}.$ (18)
In this clever covariate estimator, survey weights must be incorporated in three places: estimating the propensity score model $e_{\text{sp}}(\mathbf{X};\boldsymbol{\beta})$, the ratio $h(\mathbf{X}_{i})/p(\mathbf{X}_{i})$ for reweighting predictions, and the balancing weights $\omega_{z,p}(\mathbf{X}_{i})$ serving as clever covariates. The outcome models, including the base models $m_{z}(\mathbf{X};\boldsymbol{\alpha}_{z})$ and the refitted $m_{z}^{\text{CVR}}\left(\mathbf{X};\boldsymbol{\alpha}_{z}^{\text{CVR}}\right)$, are fitted without survey weight adjustment. The clever covariate estimator is asymptotically equivalent to the one-step moment estimator, sharing the same influence function and robustness properties. When $h(\mathbf{X})\propto 1$, the estimator achieves double robustness. When $h(\mathbf{X})$ depends on $e_{\text{sp}}(\mathbf{X})$, the estimator is singly robust with respect to the propensity score model; if the propensity score model is misspecified, the estimator converges to the shifted estimand $\widetilde{\tau}_{h}$ as discussed in Section 3.1.
@@
In the absence of survey sampling, Kang and Schafer [26] pointed out that, when the outcome models $m_{1}\left(\mathbf{X};\boldsymbol{\alpha}_{1}\right)$ and $m_{0}\left(\mathbf{X};\boldsymbol{\alpha}_{0}\right)$ are already correctly specified, the clever covariates as coarse summaries of the covariates may lead to slight overfitting by adding mean-zero noise through weighted residuals. When the survey-weighted propensity score model is correctly specified, the clever covariate method ensures consistent estimation of target parameters, provided that the conditional mean of the outcome varies smoothly with both the survey-weighted propensity score and the sampling probabilities. In theory, this smooth variation remains essential for maintaining the consistency and robustness of the clever covariate method, even in complex survey designs. However, incorporating survey weights twice in constructing the clever covariates may potentially exacerbate the extrapolation problem noted by Robins et al.[27] because the inflation of extreme value can lead to finite-sample bias, making the clever covariate estimator less stable compared to the weighted regression methods introduced in Section 3.4.